OpenStudy (anonymous):

Find cos theta if sin theta -(12/13)= and tan theta > 0

OpenStudy (xapproachesinfinity):

hmmm trying to figure out what you wrote you want to find \[\cos \theta \] given that \[\sin \theta=-\frac{12}{13}\] and \[ \tan \theta >0\]

OpenStudy (xapproachesinfinity):

so i that your question, just making sure because you wrote incorrectly

OpenStudy (xapproachesinfinity):

hey! yu there?

OpenStudy (xapproachesinfinity):

helllllllllllllllllllllllllllllllllllllllllo!!!

OpenStudy (anonymous):

sorry yea thats is

OpenStudy (anonymous):

@xapproachesinfinity ?

OpenStudy (xapproachesinfinity):

okay so since we have sin to be negative and tan positive there is only one thing and that is we are working on the third quadrant so cos must also be negative